Nation’s Capital Fitness, Inc. operates a chain of fitness centers in the Washington, D.C., area. The firm’s controller is accumulating data to be used in preparing its annual profit plan for the coming year. The cost behavior pattern of the firm’s equipment maintenance costs must be determined. The accounting staff has suggested the use of the high-low method to develop an equation, in the form of Y = a + bX, for maintenance costs. Data regarding the maintenance hours and costs for last year are as follows:

Month

Hours of Maintenance
Service

Maintenance
Costs

January

540

$

5,250

February

450

4,240

March

260

2,730

April

500

4,320

May

330

3,030

June

450

4,130

July

350

3,090

August

410

3,550

September

480

3,980

October

380

3,270

November

350

3,220

December

320

3,040

Total

4,820

$

43,850

Average

402

*

$

3,654

*

*Rounded

rev: 05_30_2018_QC_CS-128090

  1. -a. Compute the variable cost per hour and the fixed cost per hour at 640 hours of activity. (Round your answers to 2 decimal

-b. Why is the fixed cost per hour possibly misleading?

------------Because it will not change as the number of hours changes.

--------Because it will change as the number of hours changes.

2. Compute the predicted maintenance cost at 630 hours of activity.--------------

3. Using your answer to requirement (1), what is the variable component of the maintenance cost? (Round your answer to two decimal places.)

4. Using the high-low method of cost estimation, estimate the behavior of the maintenance costs incurred by Nation’s Capital Fitness, Inc. Express the cost behavior pattern in equation form. (Round coefficient of X to 2 decimal places and other answer to the nearest whole dollar amount.)
Answer #1

1. Variable Cost p u = Change in Cost / Change in Hours

= Cost Hours

1st 5250 540

Low Level     ( 4240) ( 450)

Change in $ 1010   90

Unit Variable Cost = 930 / 40 =11.22 per hour

Total Variable Cost= 540 * 11.22 = 6059

Total Fixed Cost = Total Cost - Total Variable Cost

= 5250 - 6059 = 0

Variable Cost is More than Total Cost

640 Hours Maintenance Cost = Variable Cost is not identify

Fixed Cost    is not identify

Change in Maintenance Cost but same in Hours in the Month of Feb & June

Feb's Maintenance Cost 4240

Jun's Maintenance Cost 4130

But Same 450 Hours

So Fixed cost treat as misleading Because it will not change as the number of hours changes.

2- predicted maintenance cost at 630 hours of activity (High Low Method)

640 Hours Maintenance Cost = Variable Cost = 630 x 9 = 5670

Fixed Cost 390 6,060

3- Variable Component: Y = a+Bx

A = Fixed Cost

B Variable cost per Hour

X Hours

Variable Component is B

4-

High Low Method = Cost Hours

High Level 5250 540

Low Level     ( 2730)    ( 260)

Change in $ 2520 280

Unit Variable Cost = 2520/ 280 =9.00 per hour

Total Variable Cost= 540 * 9 = 4,860

Total Fixed Cost = Total Cost - Total Variable Cost

= 5250 - 4860 = 390